Write the equation of the line with the given slope and containing the given point.

Slope - 9/10; through (-4,0)
Slope-Intercept Form______________________
Standard Form_____________________________

Answer by Cromlix(4381)

Using y - b = m(x - a)
y - 0 = 9/10 (x - (-4))
y = 9/10x + 4
Multiply through by 10
10y = 9x + 40
Hope this helps.
:-)